one number is 15 more than another. the sum of twice the larger number and 3 times the smaller number is 180. whats the numbers?
Answer by chriswen(106)

Let x be the smaller number.
Let x+15 be the larger number.
The sum of twice the larger number and 3 times the smaller number is 180.
3x+2(x+15)=180
3x+2x+30=180 distribution law
5x+30=180
5x+30-30=180-30
5x=150
5x/5=150/5
x=30
x+15=45
Therefore the numbers are 30 and 45.